OpenStudy (anonymous):

Use the graph showing Eva's account balance to answer the question that follows. a line is shown passing through the points 0 comma 525, 4 comma 556 point 5, 9 comma 595 point 88 What is the interest rate on Eva's account?

ganeshie8 (ganeshie8):

"rate" or "slope" or "gradient" all refer the same thing : how y varies with x

ganeshie8 (ganeshie8):

to get the interest rate, we can find the slope of the line. u see ?

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

since the graph is going through points \( (0, 525)\ and\ (4,556.5)\) \(\huge rate = \frac{y2-y1}{x2-x1} = \huge \frac{556.5-525}{4-0} = 7.875\)

OpenStudy (anonymous):

its multiple choice though 7.1% 6.6% 5.7% 1.5% ?

ganeshie8 (ganeshie8):

its none of the above.. we got 7.875 %

ganeshie8 (ganeshie8):

u have nay more infor... question says, some graph is shown ?

OpenStudy (anonymous):